OpenStudy (anonymous):

Find dy/dx by implicit differentiation. 4x2 + 5xy - y2 = 6

OpenStudy (anonymous):

\[4x^2+5xy-y^2=6\]

OpenStudy (anonymous):

8x + 5y + 5xdy/dx - 2ydy/dx = 0 dy/dx = (8x+5y)/(2y-5x)

OpenStudy (anonymous):

please show me step by step so that i can follow up what i am doing

OpenStudy (anonymous):

isolate the dy/dx

OpenStudy (amistre64):

aint nuthing different about implicit than normal;

OpenStudy (anonymous):

4x^2 becomes 8x naturally 5xy we need to use the product rule so 5y + 5xy' (taking dy/dx y becomes (1)y' from chain rule) -y^2 comes to be -2y * y' from chain rule

OpenStudy (amistre64):

\[4x^2 + 5xy - y^2 = 6 \] \[\frac{d(4x^2)}{dx}+\frac{d(5xy)}{dx}-\frac{y^2)}{dx}=\frac{d(6)}{dx}\]

OpenStudy (anonymous):

you need to keep in mind that you are taking the derivative with respect to x and not to y

OpenStudy (amistre64):

\[\frac{dx}{dx}8x+\frac{dx}{dx}5y+5x\frac{dy}{dx}-2y\frac{dy}{dx}=\frac{dx}{dx}0\]

OpenStudy (anonymous):

i know kanade but its gets really hard to understand some of the function

OpenStudy (amistre64):

\[{dx\over dx}=1; \frac{dy}{dx}=y'\] \[8x +5y +5x.y' -2y.y' = 0\]

OpenStudy (anonymous):

ok now it looks good..but let me see if i got the same answer

OpenStudy (amistre64):

\[5x.y' -2y.y' = -8x-5y\] \[y'(5x -2y)=-8x-5y\] \[y' = \frac{-8x-5y}{5x-2y}\]

OpenStudy (anonymous):

when you take dy/dx y^2 (with respect to x) y^2 is treated like a function but you dont know what that function is so like if you take cos^2(SOME FUNCTION) you would do 2(-sin(SOME FUNCTION))d/dx(SOME FUNCTION)

OpenStudy (amistre64):

you are used to the "baby" derivatives. \[y=x^2\] \[\frac{d(y)}{dx}=\frac{d(x^2)}{dx}\] \[1*\frac{dy}{dx}=\frac{dx}{dx}*2x\] \[y' = 2x\] but that is the same process you use in 'implicit' \[y^2 = x^2\] \[\frac{d(y^2)}{dx}=\frac{d(x^2)}{dx}\] \[2y\frac{dy}{dx}=\frac{dx}{dx}2x\] \[2y.y' = 2x\] \[y' = \frac{2x}{2y}=\frac{x}{y}\]

OpenStudy (anonymous):

ok cool thanks soo much